# Generating Fock-state superposition from coherent state by quantum measurement

### v1

@@ V Conclusion
In summary, we design a protocol based on a sequence of projective measurements on the ancillary atomic system to generate desired Fock states and superposed Fock states from a coherent state of a resonator. The measurements are separated by dozens of joint free evolutions of the resonator and the ancillary system that are coupled with a simple JC interaction. Through the analysis over the population reduction factors induced by the effective time-evolution operator for the target resonator, we can find optimal strategies with varying duration of the evolution-and-measurement cycles. Our protocol can therefore utilize less than $30$ measurements to create a Fock state $|n\rangle$ or a superposed Fock state $(|0\rangle+|n\rangle)/\sqrt{2}$ of $n\sim 10$ with a close-to-unit fidelity. The success probability of our conditional protocol is exclusively determined by the overlap between the initial state and the target state of the resonator. Moreover, the protocol can be generalized to prepare a multi-excitation Bell state in a double-resonator system assisted by a qutrit. Our work thus proves quantum measurement to be a powerful tool to manipulate composite systems and presents a path to the generation of high-level Fock states and entangled states.
@@ Acknowledgments
We acknowledge grant support from the National Natural Science Foundation of China (Grant No. 11974311).

### v2

@@ V Conclusion
In summary, we have constructed a protocol based on a sequence of projective measurements on the ancillary atomic system to generate Fock states and superposed Fock states from a coherent state of a resonator. The measurements are separated by dozens of rounds of joint free evolutions of the resonator and the ancillary system, which are coupled with a JC interaction. Through analysis of the population-reduction factors induced by the effective time-evolution operator for the target resonator, we found optimal strategies with varying duration of the evolution-and-measurement cycles. In the absence of decoherence, our protocol can therefore use fewer than $30$ measurements to create a Fock state $|n\rangle$ or a superposed Fock state $(|0\rangle+|n\rangle)/\sqrt{2}$ of $n\sim 10$ with a close-to-unit fidelity. Moreover, it can be generalized to prepare a Bell-like state $(|00\rangle+|nn\rangle)/\sqrt{2}$ in a double-resonator system assisted by a qutrit. Our protocol does not require any elaborate parametric control and external driving on the system. The success probability of our conditional protocol is determined by the overlap between the initial state and the target state of the resonator. In the presence of decoherence, the attainable fidelity for a given target state is limited by the decay to the lower states and an optimized number of measurements emerges for generating a Bell-like state. Our work thus proves quantum measurement to be a powerful tool to manipulate composite systems and presents a path to the generation of high-level Fock states and entangled states.
@@ Acknowledgments
We acknowledge financial support from the National Natural Science Foundation of China (Grant No. 11974311).